# Unveiling Attacks Ahead of Exploit: adversarial Contract Detection Based on DeFi Attack Features

### 14 Jan 2024 v1

@@ 5.3 Feature Importance
To assess the impact of different feature categories, we divided the features into two major classes: one focused on transaction-related attributes (Creator Attribute, Transaction Data, Verification Status), and the other related to the contract’s code itself (Implemented Function, Function Call, Contract Type). We conducted tests using only one feature category at a time, and the results are presented in Table 3.
@@
Clearly, when compared to using a combination of features, training classifiers with a single category of features leads to a certain degree of performance degradation for all models. This indicates that each category of features plays a valuable role in the detection of adversarial contracts. Additionally, Figure 5 illustrates the feature importance of all features for the LightGBM model.
Figure 5: Top 15 LightGBM Feature Importance
6 Limitations and Future Work
@@
Detection Scope. Our current method has certain limitations in the scope of detection. In terms of feature selection, our emphasis is DeFi attacks involving financial losses, specifically tailored to common attack scenarios. Consequently, we may miss attacks that do not result in financial losses and certain specialized attack types. For instance, during evaluation, we found that a small number of undetected attacks related to NFTs, for which we did not have specifically designed features. Moreover, we have observed that some DeFi attackers, similar to front-running bots, encapsulate function calls related to attacks within calldata. As the attacking logic employing this approach is not included in the contract, it also poses difficulties for detection.
@@
Performance. The current feature extraction process relies on RPC providers and blockchain explorer APIs. While these components can be replaced by running our custom nodes, there is currently a lack of robust open-source decompilation tools. Solely relying on Dedaub’s online decompilation service results in slow response times, making real-time detection impractical.
In the future, we consider adjusting certain features to accommodate evolving attack types. Furthermore, there is room for constructing our own decompilation module. This would not only improve feature extraction performance but also allow for the extraction of more comprehensive features from bytecode.
7 Related Work
In this section, we discuss existing works that are related to our work. In particular, we first discuss existing works for adversarial transaction detection, then present the papers related to abnormal contract detection.
7.1 Adversarial Transaction Detection
@@
Some researches[38, 50, 31, 57] modify EVM to obtain more detailed transaction execution information and provide scalable interfaces for online detection of various types of adversarial transactions. Sereum[50] uses taint analysis to monitor runtime data flows during the execution of transactions to detect reentrancy attacks. EthScope[57] is a scalable Ethereum attack detection framework which stores historical data into an cluster database and adds dynamic taint analysis to the client to analyze transactions.
@@
On the other hand, some researches[60, 63, 54, 58, 46] achieve adversarial transaction detection through forensic analysis without modifying EVM. TxSpector[60] is a logic-driven framework designed to analyze Ethereum transactions, identifying attacks according to user-defined detection rules by encoding the transaction trace into logical relationships. Qin et al.[46] propose an Execution Property Graph and a graph traversal technique which can swiftly identify potential adversarial transactions efficiently.
7.2 Abnormal Contract Detection
@@
Forta‘s [21] method combines Natural Language Processing (NLP) techniques to identify adversarial contracts at the bytecode level, and apart from it, there is no existing work that specifically focuses on the detection of adversarial contracts. The majority of abnormal contract detection focuses on Ponzi schemes and honeypots. They share similarities in methods and can be categorized into code similarity detection[27, 52], machine learning[61, 40, 41, 39, 29, 33], and symbolic execution techniques[32, 53].
Ponzi scheme. Ponzi scheme is a financial fraud that attracts users with promises of high profits. With the emergence of blockchain technology, it has provided many opportunities for scammers. Zheng et al.[32] constructed a larger dataset of Ponzi scheme than previous work[33] and proposed a Multi-view Cascade Ensemble method for detecting smart Ponzi schemes on Ethereum.
@@
Honeypots. Honeypots are a distinct type of contracts designed to entice victims into traps, employing seemingly vulnerable structures that conceal hidden risks rather than relying on traditional deceptive methods. Torres et al.[53] developed a taxonomy of honeypot techniques and built a tool called HONEYBAGER that employs symbolic execution and well defined heuristics to expose honeypots.
8 Conclusion
@@
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for adversarial contracts. Based on DeFi attack patterns, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones. The evaluation show that our method performs exceptionally well. The F1-Score for LightGBM-based classifier is 0.9434, with a remarkably low false positive rate of only 0.12

### 02 Feb 2024 v2

@@ 5.3. Feature Importance
To assess the impact of different feature categories, we divided the features into two major classes: one focused on the contract metadata, including Creator Attribute, Transaction Data, and Verification Status, the other related to the contract functionality, including Implemented Function, Function Call, and Contract Type. We conducted tests using only one feature class at a time, and the results are presented in Table 6.
@@
Clearly, when compared to using all features, training classifiers with a single class of features leads to a certain degree of performance degradation for all models. This indicates that each category of features plays a valuable role in the detection of adversarial contracts. Moreover, it’s worth noting that the performance varies across different models when utilizing the two feature classes separately. Random Forest, MLP, and LightGBM models tend to perform better when using functionality features, while the other two models exhibit better performance with metadata features. This observation suggests that the importance of different features varies depending on the model used.
Figure 11 illustrates the top 15 important features of the lightGBM-based classifier when all features are used. It is evident that various feature categories in this classifier contribute significantly to its performance.
5.4. Failure Analysis
In this section, we analyze the failure instances of the LightGBM-based classifier, where the classifier made incorrect identifications during testing. Specifically, there is one false negative and four false positives. Table 8 presents these failure contracts along with a subset of their important features.
5.4.1. False Negative
The adversarial contract with address 0x0f41f9 stands out as the sole false negative instance, erroneously classified as benign by the classifier. While it exhibits some characteristics typically associated with adversarial contracts, such as an Anonymous fund source, lack of verification, a single public function, and relatively low input data and gas consumption, it notably lacks the implementation of the flashloan callback and does not involve token-related calls. These deviations from the typical adversarial contract might have contributed to the misclassification by the classifier.
This incident occurred on September 29, 2023, resulting in approximately 4,000 USD in losses. Upon analyzing the victim contract and the adversarial transaction, it becomes apparent that the vulnerability exploited in this attack was an access control flaw that led to arbitrary external call. In the adversarial contract, the attack was completed with only three external calls. (1) It first invoked the update function within the victim contract. This function could modify critical variables but lack proper access control, enabling anyone to call it. (2) Next, the adversarial contract exploited the functionCallWithValue function of the victim contract, transferring its native tokens to adversarial contract specified in the previously called update function. (3) Finally, the acquired tokens were transferred to the attacker. This attack was relatively simple, with losses directly in native tokens and no involvement of token-related calls, suggesting that our detection method may have limitations when confronted with adversarial contracts that employ straightforward attack flow.
5.4.2. False Positive
The four false positive instances share certain characteristics resembling typical adversarial contracts, such as non-Safe fund source labels, unverified status, the usage of token-related calls, and the ability to self-destruct. These particular features might have led the classifier to identify them as adversarial contracts.
Arbitrage. The contract 0xdac3d4 was found to be involved in arbitrage activities through an analysis of its transaction history. Arbitrage involves profiting from price differences in tokens across different platforms. Specifically, this contract was used for the NFT arbitrage. Taking its first transaction after deployment, transaction 0x25637d, as an example, the contract initially purchased an NFT on the Blur (blur, 9) for 0.3 Ether and immediately sold it on the Sudoswap (sudoswap, 56) for approximately 0.314 Ether, thereby profiting from the price difference. The entire process was executed within a single transaction using the contract and leveraged the private service of flashbots to mitigate the risk of failure.
SniperBots. The contract 0x1d2cbc has already self-destructed. Upon analyzing its transaction history, it was observed that nearly all of its transactions involved purchasing ERC-20 tokens with Ether. Furthermore, the purchase transaction and the transaction for creating the liquidity pool corresponding to the target token both occurred within the same block. Consequently, this contract was actually used by sniper bots (cernera2023token, 14). These bots engage in buying tokens on liquidity pools as soon as they are listed, allowing them to ”snipe” new tokens at their initial price before anyone else, with the intention of selling them at a profit once their value increases. The contract 0x70d2f0 was also utilized by similar sniper bots.
Phishing. The contract 0xfd594e has relatively simple code. An analysis of its transaction history reveals a consistent pattern in all transactions, which is phishing activity in fact. Specifically, it was employed for zero-value transfer phishing. In this phishing method, scammers begin by generating fake addresses with characters matching the first and last few characters of the victim’s recent interaction address. Subsequently, they employ the transferFrom function to perform zero-value transfer to the victim’s address. This action results in the appearance of the fake address in the victim’s transaction history, misleading users with the intent of causing them to inadvertently send tokens to this fraudulent address, thereby stealing the victim’s tokens.
These false positives suggest that contracts used for arbitrage, as well as those employed in sniper bots, phishing, and other suspicious activities aimed at financial gain, bear certain similarities to adversarial contracts involved in DeFi attacks. These may represent the primary contract types responsible for the false positives in our detection methodology.
6. Limitations and Future Work
@@
Detection Scope. Our current method has certain limitations in the scope of detection. In terms of feature selection, we rely on existing attack patterns, with a specific focus on DeFi attacks involving financial losses, tailored to prevalent attack scenarios. Consequently, we may miss attacks that do not result in financial losses and certain adversarial contracts with novel attack vectors. Furthermore, in the cat-and-mouse game of DeFi attack and defense, we have observed that some DeFi attackers, similar to front-running bots, encapsulate function calls related to attacks within the input data of adversarial transactions. As the attacking logic employing this approach is not included in the contract, it also poses difficulties for detection.
@@
Performance. As analyzed in section §5.4, our classifier struggles to correctly categorize adversarial contracts with overly simplistic attacking logic, as well as contracts designed for suspicious activities aimed at financial gain, which significantly impacts the accuracy of our detection system. Additionally, the current feature extraction process relies on RPC providers and blockchain explorer APIs. While these components can be replaced by running our custom nodes, as of writing, there is a lack of robust open-source decompilation tools. Solely relying on Dedaub’s online decompilation service results in slow response times, making real-time detection impractical.
In the future, we will continue to refine our contract dataset, which involves not only supplementing the adversarial contracts but also gathering a more diverse range of benign contracts from a broader time span. Additionally, we plan to delve deeper into the nuanced differences between contracts utilized in suspicious activities such as arbitrage, sniper bots, phishing, and contracts implicated in DeFi attacks, helping us extract more representative features to reduce false positives while maintaining efficient detection capability for adversarial contracts. We are also contemplating a more granular classification of adversarial contracts based on specific attacking logic they employ, enabling us to identify the vulnerabilities that the attackers intent to exploit in the victim contract. Furthermore, we recognize the potential benefits of developing our own decompilation module. This would not only improve the real-time performance of feature extraction but also allow for more comprehensive features from bytecode.
7. Related Work
In this section, we introduce two categories of closely-related work.
7.1. Adversarial Transaction Detection
@@
Some researches (chen2020soda, 17, 38, 53, 69) modify EVM to obtain more detailed transaction execution information and provide scalable interfaces for online detection of various types of adversarial transactions. Sereum (rodler2018sereum, 53) uses taint analysis to monitor runtime data flows during the execution of transactions to detect reentrancy attacks. SODA (chen2020soda, 17) is a generic online detection framework that can be used to detect various adversarial transactions based on predefined rules. EthScope (wu2022time, 69) is a scalable Ethereum attack detection framework which stores historical data into an cluster database and incorporates dynamic taint analysis into the client to analyze transactions.
@@
On the other hand, some researches (gai2023blockchain, 37, 49, 64, 70, 72, 76) achieve adversarial transaction detection through forensic analysis. TxSpector (zhang2020txspector, 72) is a logic-driven framework designed to analyze Ethereum transactions, identifying attacks according to user-defined detection rules by encoding the transaction trace into logical relationships. DeFiScanner (wang2022defiscanner, 64) is the first deep-learning-based attack detection system, capable of identifying adversarial transactions based on features from both the account and transaction levels. BLOCKGPT (gai2023blockchain, 37) applies large language model into abnormal transaction detection and demonstrates strong performance in detecting adversarial transactions. Qin et al. (qin2023towards, 49) propose an Execution Property Graph and a graph traversal technique which can swiftly identify potential adversarial transactions efficiently.
7.2. Abnormal Contract Detection
@@
Forta‘s (FortaNetwork, 35) method combines Natural Language Processing (NLP) techniques to identify adversarial contracts at the bytecode level, and apart from it, there is no existing work that specifically focuses on the detection of adversarial contracts. The majority of abnormal contract detection focuses on Ponzi schemes and honeypots. They share similarities in methods and can be categorized into code similarity detection (bartoletti2020dissecting, 7, 57), machine learning (camino2020data, 12, 18, 19, 39, 40, 41, 74), and symbolic execution techniques (chen2021sadponzi, 20, 61).
Ponzi scheme. The traditional Ponzi scheme is a financial fraud that promises high returns to early investors by using funds from new investors, ultimately collapsing when there are not enough new investors, leaving most participants with losses. The emergence of smart contracts has enabled the replication of Ponzi schemes on blockchain platforms. Chen et al. (chen2018detecting, 19) focused on extracting features based on the contract bytecode and the transaction information of the contract. Ji et al. (ji2023ponzi, 41) extracted features based on the control flow graph of the contract. Zheng et al. (zheng2023securing, 74) constructed a larger dataset of Ponzi scheme than previous work (chen2018detecting, 19) and proposed a Multi-view Cascade Ensemble method without relying on contract interaction information.
@@
Honeypots. Honeypots are a distinct type of contracts designed to entice victims into traps, employing seemingly vulnerable structures to conceal risks rather than relying on traditional deceptive methods. Torres et al. (torres2019art, 61) developed a taxonomy of honeypot techniques and built a tool called HoneyBadger that employs symbolic execution and well defined heuristics to expose honeypots. Subsequently, Camino et al. (camino2020data, 12) introduced a data science approach based on the transaction behavior of contracts. Chen et al. (chen2020honeypot, 18) constructed a classification model based on the opcode of contracts.
8. Conclusion
@@
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for adversarial contracts. Based on this dataset, we found that the proportion of adversarial transactions sent using privacy services has exhibited a significant upward trend, thereby evading traditional adversarial transaction detection methods. Through the analysis of DeFi attack patterns, we extract six categories of features across multiple dimensions, such as attributes of the contract creator, functions implemented in the contract, and types of contracts. Utilizing these features, we build a machine learning classifier that can effectively distinguish adversarial contracts from benign ones. The evaluation show that our method performs exceptionally well. The F1-Score for LightGBM-based classifier is 0.9541, with a remarkably low false positive rate of only 0.15%. By analyzing cases where the classifier made incorrect identifications, we point out contracts that could be confused with adversarial contracts used in DeFi attacks.